Use a sum or difference formula to find the exact value of the following.
cos(5pi/7)cos(13pi/28) + sin(5pi/7)sin(13pi/28) =

Answer by lwsshak3(11628)

Use a sum or difference formula to find the exact value of the following.
cos(5pi/7)cos(13pi/28) + sin(5pi/7)sin(13pi/28)
***
Identity: cos(x-y)=cosx*cosy+sinx*siny
cos(5pi/7)cos(13pi/28) + sin(5pi/7)sin(13pi/28)=cos(5π/7-13π/8)=cos(40π/56-91π/56)=cos(-51π/56)